Ralph advertises to sell cookies for Php 200 a dozen. He

sells 75 dozen and decides that he can charge more. He raises the price

to Php 300 a dozen and sells 45 dozen. What is the elasticity of demand?


1
Expert's answer
2021-12-03T13:08:38-0500

 The elasticity of demand (PED) is:


"PED=\frac{\Delta Q}{\Delta P}"


where;

"\Delta Q=\frac{Q_2-Q_1}{Q_1}=\frac{45-75}{75}=-0.4"


"\Delta P=\frac{P_2-P_1}{P_1}=\frac{300-200}{200}=0.5"


"PED=\frac{-0.4}{0.5}=-0.8"


 The elasticity of demand is -0.8 that is negative. This means the increase in price leads to decrease in quantity demanded in the market.